Electrostatics-Firstname Lastname 13: A: An iron nucleus (Q-4.16x101 C and M- 9.35x1020 kgis initially moving horizontally with a speed Vo-2x10 m/s. Upon reaching pointA, it enters a region with a constant vertical electric field of 8000N/C. Which of the two paths does the charge follow? Explai B: Determine the kinetic energy of the particle 3x10 s later. Hint: the motion of this charge is qualitatively identical to projectile motion.
Answer #1

A) since iron nucleus is a positive charge

force on it will act in the direction of electric field

so it will follow path A

B) F = qE

a = qE/ m = 4.16*10^-18*8000/9.35*10^-26

= 3.56*10^11 m/s2

v = u + at

v = 2*10^6 + 3.56*10^11*3*10^-6

v = 3.067*10^6 m/s

KE = 0.5mv^2 = 0.5*9.35*10^-26*(3.067*10^6)^2

= 4.4*10^-13 J